# Computational Copyright: Towards A Royalty Model for AI Music Generation Platforms

### v1

@@ Abstract
The advancement of generative AI has given rise to pressing copyright challenges, particularly in music industry. This paper focuses on the economic aspects of these challenges, emphasizing that the economic impact constitutes a central issue in the copyright arena. The complexity of the black-box generative AI technologies not only suggests but necessitates algorithmic solutions. However, such solutions have been largely missing, leading to regulatory challenges in this landscape. We aim to bridge the gap in current approaches by proposing potential royalty models for revenue sharing on AI music generation platforms. Our methodology involves a detailed analysis of existing royalty models in platforms like Spotify and YouTube, and adapting these to the unique context of AI-generated music. A significant challenge we address is the attribution of AI-generated music to influential copyrighted content in the training data. To this end, we present algorithmic solutions employing data attribution techniques. Our experimental results verify the effectiveness of these solutions. This research represents a pioneering effort in integrating technical advancements with economic and legal considerations in the field of generative AI, offering a computational copyright solution for the challenges posed by the opaque nature of AI technologies.
@@ 1 Introduction
Recent advancements in generative AI have significantly impacted creative industries, leading to a surge in AI-generated content across art, music, literature, and software. This rapid evolution has raised complex legal challenges, especially concerning copyright issues (Henderson et al., 2023; Samuelson, 2023; Sag, 2023; Franceschelli and Musolesi, 2022). Copyright laws cover a range of rights, including protection of original works, controlling their reproduction, and managing the distribution of profits from these works. The emergence of generative AI poses multifaceted challenges in this regard, as it blurs the lines of authorship and originality.
@@
Arguably, central to these challenges is the economic impact, particularly evident in the music industry. The widespread availability of music on platforms like Spotify and YouTube, where copyright owners are compensated through royalties, illustrates this point. This practice not only demonstrates that economic incentives are a primary reason for making music publicly accessible, but also highlights the centrality of economic rights in copyright protections. This trend is reflective of a broader truth: economic considerations are at the heart of the U.S. copyright law, where a primary goal is to stimulate creativity by ensuring that creators are adequately compensated. There has also been ongoing debate about whether training generative AI with copyrighted content aligns with the fair use doctrine [^1] See Section 107 of the Copyright Act: https://www.copyright.gov/title17/92chap1.html#107.. However, it is increasingly argued that fair use may not apply if the AI generated content competes with the original market for the data (Henderson et al., 2023). These issues underscore the economic impact as a crucial aspect of copyright challenges in generative AI.
@@
However, there has been little to no technical solutions effectively solving the aforementioned challenge. Existing efforts have focused on preventing generative AI from generating content similar to its training data (Vyas et al., 2023; Chu et al., 2023; Li et al., 2023). This approach, while helpful, may not fully address the broader economic implications of AI-generated content. Addressing the economic aspect of the copyright challenges is challenging as it requires a solution that integrates technical advancement into business agreements.
@@
The challenge is also pressing. Without effective technical solutions for proper royalty distribution, regulatory bodies are faced with a dilemma between stifling innovation and compromising the interests of copyright owners. As it stands, numerous AI music generation platforms are navigating these uncharted waters, operating in legal gray areas and leaving the rights of copyright owners inadequately protected (Drott, 2021; Clancy, 2021).
@@
This paper aims to bridge this crucial gap by proposing potential royalty models for revenue sharing from AI music generation platforms. Specifically, we design the royalty model by addressing the following key questions:
  1. 1.
    Who are the stakeholders?
  2. 2.
    What are the sources of revenue?
  3. 3.
    How to determine the royalty distribution for revenue sharing?
@@
The proposed approach represents an early effort to address the complex intersection of technological innovation and economic considerations in copyright law for generative AI. The complexity of the black-box generative AI technologies necessitates a computational copyright solution. This paper showcases a promising prototype towards this goal.
@@ 2 Exploring Digital Music Royalty Models Through Case Studies
In this section, we examine the royalty models in the digital music industry through a couple of case studies.
2.1 Preliminary: The Concepts of Music Royalties

### v3

# Computational Copyright: Towards A Royalty Model for Music Generative AI

@@ Abstract
The advancement of generative AI has given rise to pressing copyright challenges, especially within the music industry. This paper focuses on the economic aspects of these challenges, emphasizing that the economic impact constitutes a central issue in the copyright arena. Furthermore, the complexity of the black-box generative AI technologies not only suggests but necessitates algorithmic solutions. Yet, such solutions have been largely missing, exacerbating regulatory hurdles in this landscape. We seek to address this gap by proposing viable royalty models for revenue sharing on AI music generation platforms. We start by examining existing royalty models utilized by platforms like Spotify and YouTube, and then discuss how to adapt them to the unique context of AI-generated music. A significant challenge emerging from this adaptation is the attribution of AI-generated music to influential copyrighted content in the training data. To this end, we present algorithmic solutions employing data attribution techniques. We also conduct a range of experiments to verify the effectiveness and robustness of these solutions. This research is one of the early attempts to integrate technical advancements with economic and legal considerations in the field of music generative AI, offering a computational copyright solution for the challenges posed by the opaque nature of AI technologies.
@@ 1 Introduction
Recent advancements in generative AI have significantly impacted creative industries, leading to a surge in AI-generated content across art, music, literature, and software. This rapid evolution has raised complex legal challenges, especially concerning copyright issues Henderson et al. (2023); Samuelson (2023); Sag (2023); Franceschelli and Musolesi (2022). A notable instance of these challenges is the recent lawsuit filed by New York Times against Microsoft and OpenAI NYT (2023). Copyright laws cover a range of rights, including protection of original works, controlling their reproduction, and managing the distribution of profits from these works. The emergence of generative AI poses multifaceted challenges in this regard, as it blurs the lines of authorship and originality.
@@
Arguably, central to these challenges is the economic impact. Taking the music industry as an example, a vast collection of music has been publicly available on platforms like Spotify and YouTube, where copyright owners are compensated through royalties. This practice not only suggests that economic incentives are a primary reason for making music publicly accessible, but also highlights the centrality of economic rights in copyright protections. This trend is reflective of a broader truth: economic considerations are at the heart of the U.S. copyright law, where a primary goal is to stimulate creativity by ensuring that creators are adequately compensated. There has also been ongoing debate about whether training generative AI with copyrighted content aligns with the fair use doctrine [^1] See Section 107 of the Copyright Act: https://www.copyright.gov/title17/92chap1.html#107.. However, it is increasingly argued that fair use may not apply if the AI generated content competes with the original market for the data Henderson et al. (2023). These issues underscore the economic impact as a crucial aspect of copyright challenges in generative AI.
@@
However, effective technical solutions addressing the aforementioned challenge have been limited or nonexistent. Existing efforts have focused on preventing generative AI from generating content similar to its training data Vyas et al. (2023); Chu et al. (2023); Li et al. (2023). This approach, while helpful, may not fully address the broader economic implications of AI-generated content. Addressing the economic aspect of the copyright challenges is challenging as it requires a solution that integrates technical advancement into business agreements.
@@
The challenge is also pressing. Without effective technical solutions for proper royalty distribution, regulatory bodies are faced with a dilemma between stifling innovation and compromising the interests of copyright owners. As it stands, numerous AI music generation platforms are navigating these uncharted waters, operating in legal gray areas and leaving the rights of copyright owners inadequately protected Drott (2021); Clancy (2021).
@@
This paper aims to bridge this crucial gap by proposing potential royalty models for revenue sharing from AI music generation platforms. Specifically, we design the royalty model by addressing the following key questions: 1) Who are the stakeholders? 2) What are the sources of revenue? 3) How to determine the royalty distribution for revenue sharing?
@@
The proposed approach represents an early effort to navigate the complex intersection of technological innovation and economic considerations in copyright law for generative AI. The complexity of the black-box generative AI technologies necessitates a computational copyright solution. This paper showcases a promising prototype towards this goal.
@@ 2 Exploring Digital Music Royalty Models Through Case Studies
In this section, we examine the royalty models in the digital music industry through a couple of case studies. Please refer to Appendix A for fundamental concepts and a few major types of music royalties that are prevalent in the industry. In order to understand the intricacies of the implementation of royalty models and their applicability to AI music generation platforms, we delve into case studies of two major platforms: Spotify and YouTube. Spotify is the largest music streaming platform in the world while YouTube is the largest video sharing platform. Both platforms have a significant amount of music content and generate revenue through multiple sources. Furthermore, despite various existing criticisms on these royalty models Marshall (2015); Trendacosta (2020), they represent the status quo of how the current digital music market works. Therefore, designing royalty models for AI music generation platforms mimicking the ones for Spotify and YouTube would be a reasonable initial step in this area. In the following sections, we will examine the royalty models of these two platforms in detail.
2.1 Case Study: Spotify
Stakeholders.
Spotify’s royalty model involves several key groups of stakeholders, in addition to the streaming platform itself. These groups [^2] Please refer to Appendix B for detailed description of these groups of stakeholders. are (1) artists and creators, (2) record labels and music publishers, (3) music rights societies and collecting agencies, (4) listeners and subscribers, and (5) advertisers.
Stakeholders in groups 1, 2, and 3 receive revenue shares from Spotify, while groups 4 and 5 contribute to the generation of Spotify’s revenue. Typically, Spotify directly interacts with stakeholders in groups 2 and 3. Individual artists and creators often have contracts with these labels, publishers, or music rights agencies, and do not directly engage with Spotify in the financial aspect of their music streaming.
Revenue Sources.
The major revenue sources of Spotify can be divided into two categories: subscription and advertisement. In 2021, premium subscriptions accounted for 88% of Spotify’s revenue while advertisements accounted for the remaining 12% Johnston (2023). The two revenue sources lead to the formation of separate revenue pools, which are also calculated separately for different countries or regions.
Royalty Distribution.
Spotify employs a straightforward pro rata model to calculate the royalty distribution for each revenue pool. The royalty for each artist or track is calculated by applying their stream share to each revenue pool. This method ensures that royalty distribution is directly proportional to the popularity and streaming frequency of each artist’s or track’s work on the platform.
2.2 Case Study: YouTube
YouTube’s model for compensating music copyright owners is multifaceted, offering various methods for monetizing the content: (1) YouTube Partner Program: Music copyright owners can join the YouTube Partner Program, uploading music (videos) to their official channels. Revenue is shared based on user views of their content; (2) Content ID: Owners can earn from videos using their music through the Content ID system. This system uses fingerprinting and machine learning to identify copyrighted content in uploaded videos and allocates revenue from these videos to the copyright owners; (3) Licensing: Owners can also license their music directly to a YouTube video for a one-time payment.
The first method resembles Spotify’s royalty model. The second and the third methods are different as they involve a third-party video creator.
Stakeholders.
Revenue Sources.
Royalty Distribution.
A crucial aspect of YouTube’s royalty model is the challenge of attributing copyright ownership in the videos. Unlike Spotify, where content attribution is straightforward through stream counts, the incorporation of copyrighted music in the user-generated videos makes this task technically demanding. The Content ID system serves as the technical foundation that enables the second and third methods of YouTube’s revenue sharing. In the second method, it identifies music in videos and allocates revenue to the copyright owners. In the third method, while synchronization licenses might be obtained through third-party licensing platforms outside of YouTube, the presence of the Content ID system encourages them to secure these licenses. Although the system has its share of flaws and has faced criticism Van der Sar (2021); McKay (2011); Trendacosta (2020); Saadatpanah et al. (2020), including issues with false positives and negatives, it is still broadly embraced by the music industry.
A notable pattern of the royalty models on digital music/video platforms is that the payment is not directly made to each individual piece of music work [^3] The licensing model on YouTube, where video creators directly buy synchronization licenses for individual music, is an exception. However, note that this model is also enabled by the existence of the Content ID system.. Rather, the platforms typically follow two main steps: 1) formation of revenue pools and 2) distribution based on access frequency.
3 Potential Royalty Models for AI Music Generation Platforms
3.1 The Business Models
While the landscape of AI music generation is still rapidly evolving, there have been a few common business models emerging McFarland (2023). We summarize these business models in terms of services, stakeholders, and revenue sources [^4] Please refer to Appendix C for more detailed AI music generation platforms business models..
The backbone of AI music generation platforms is generative AI trained on a large corpus of existing music, which often includes copyrighted music. With the generative AI, the platforms offer a variety of services to meet different needs of end users. The potential stakeholders involved in AI music generation platforms have significant overlaps with those on traditional music platforms, as summarized in the five groups in Section 2.1. The platforms have several different ways for generating revenues, such as subscription fees, licensing, advertisements and costom composition fees.
3.2 Potential Royalty Model Designs
Particularly, the business models for AI music generation platforms align with the pattern identified in the case study, where the revenue is channeled through the platform, rather than directly compensating for each individual piece of music.
Formation of Revenue Pools.
Similar to Spotify and YouTube, these platforms would first accumulate revenue, forming distinct pools based on different criteria such as revenue sources (subscriptions, advertisements, licensing fees).
Distribution Based on Access Frequency.
The revenue from each pool would then be distributed based on the frequency at which each copyrighted work included in the training corpus is accessed during the service.
4 Attributing AI-Generated Music to Copyrighted Content
4.1 Formulation of the Attribution Problem
Attributing the influence of copyrighted training content on generated music essentially asks the question: “To what extent does each piece of music in the training corpus influence a specific piece of AI-generated music?” The quantifiable definition of “influence” can potentially be subjective. We suggest two perspectives to define the “influence”, one inspired by the machine learning literature and the other comes from the domain knowledge of music.
Perspective 1: Data Attribution.
The data attribution problem in machine learning refers to the task of identifying and quantifying the contribution of each data point in the training dataset to a given output of a machine learning model. Formally, this problem is often framed as follows: How does the removal of a particular data point from the training dataset and subsequent retraining of the model affect its output? This change in output serves as a measure of the removed data point’s influence on that specific model output Koh and Liang (2017). In the context of AI music generation, we can define the influence of a piece of training music on a piece of generated music in terms of the change in the likelihood of the model producing that generated music, assuming the model is retrained after removing the training music piece.
Perspective 2: Musical Influence.
4.2 Data Attribution for Music Generation
In this section, we first introduce formal definitions of AI music generation and the data attribution problem for music generation. Subsequently, we propose an algorithmic solution to quantitatively estimate the influence of each training piece on a particular piece of generated music.
4.2.1 AI Music Generation
In the field of AI music generation, there are two major paradigms: waveform music generation and symbolic music generation Manzelli et al. (2018). Waveform music generation involves the direct synthesis of a music’s waveform, with examples including WaveNet Oord et al. (2016). Symbolic music generation involves creating music in a symbolic format, such as the Musical Instrument Digital Interface (MIDI) format. This paper focuses on symbolic music generation.
Symbolic Music Generation.
We start by introducing the notations and the formal definitions of symbolic music generation. Symbolic music is characterized as a series of discrete events that collectively represent a segment of music. A music segment $m$ with $k$ events can be expressed as $(e_{1},e_{2},\ldots,e_{k})$. Here, each $e_{k}\in\mathcal{V}$ represents a specific event from a vocabulary set $\mathcal{V}$ that defines all the valid symbolic events.
A symbolic music generation model, denoted as $h$, takes a prompt music segment and calculates the subsequent event’s conditional probability distribution $P(e_{k}\mid e_{k-p},e_{k-p+1},\ldots,e_{k-1};h)$, where $p$ indicates the length of prompt. Suppose the size of the vocabulary set is $|\mathcal{V}|=V$. The model $h$ can be represented as a classification neural network with $V$ output units correspond to a probability distribution over the vocabulary set. This type of model formulation is known as autoregressive model, which is one of the most popular model family for symbolic music.
4.2.2 Data Attribution for Symbolic Music Generation
Attribution Scores.
$I(m_{i},\hat{m})=f(\hat{m},h_{S\setminus\{m_{i}\}})-f(\hat{m},h_{S}),$ (1)
Two Levels of Attribution.
Estimating the Attribution Scores.
Directly calculating $I(m_{i},\hat{m})$ requires retraining a model for for each training data point $m_{i}$, which is computationally prohibitive. Fortunately, there has been a rich literature on efficient data attribution methods Hammoudeh and Lowd (2022), primarily designed for classification models. Furthermore, we shall see that these methods can be easily adapted to the autoregressive generative models like Music Transformer.
In particular, since generating one event can be viewed as a classification problem, we can directly apply existing data attribution methods for classification models to event-level attribution. For the segment-level attribution, when the utility function is defined as the log-likelihood, i.e., $f(\hat{m},h)=\log P(\hat{m};h)$. We observe that, assuming $\hat{m}=(e_{k},\ldots,e_{l})$ for some $k<l$, then by Bayes’ rule, $\log P(\hat{m};h)=\Sigma_{i=k}^{l}f(\{e_{i}\},h).$
$I(m_{i},\hat{m})$ $=f(\hat{m},h_{S\setminus\{m_{i}\}})-f(\hat{m},h_{S})$
$=\Sigma_{i=k}^{l}\left(f(\{e_{i}\},h_{S\setminus\{m_{i}\}})-f(\{e_{i}\},h_{S})\right)$
$=\Sigma_{i=k}^{l}I(m_{i},\{e_{i}\}),$
Several off-the-shelf data attribution methods can be applied to estimate the attribution scores for the autoregressive symbolic music generation models. We denote the estimated attribution score for $m_{i}$ on $\hat{m}$ as $\hat{I}(m_{i},\hat{m})$.
4.3 Experimental Evaluation
We conduct experiments that extend two data attribution methods, TracIN Pruthi et al. (2020) and TRAK Park et al. (2023), for the Music Transformer model Huang et al. (2018), evaluated on the MAESTRO dataset Hawthorne et al. (2019) [^5] Please refer to Appendix D for detailed experimental setup..
4.3.1 Evaluation Metrics
We introduce evaluation metrics formalizing the two perspectives of influence in Section 4.1.
Retraining Rank Correlation.
To evaluate the estimated attribution scores from the data attribution perspective, we can compare them with the ground truth attribution scores defined in Eq. (1). In the data attribution literature Koh and Liang (2017); Ilyas et al. (2022), the comparison is typically measured by Spearman’s rank correlation. Formally, for a training dataset with $n$ data points, one will calculate the rank correlation between $\{\hat{I}(m_{i},\hat{m})\}_{i=1}^{n}$ and $\{I(m_{i},\hat{m})\}_{i=1}^{n}$.
Musical Similarity.
For the musical influence perspective, there are multiple aspects mentioned in Section 4.1. In our study, we focus on the similarity of musical styles. A common approach to quantitatively evaluate musical style similarity is by extracting features from the music Slaney et al. (2008). In this study, we identify three features used in Spotify API [^6] https://developer.spotify.com/documentation/web-api/ to characterize a piece of music. Loudness measures the overall velocity of a music segment. We define it as the average velocity of events within the segment. Key measures the average pitch height of all events in the music segment. Duration measures the total length of the music segment in time, calculated as the sum of the time deltas of all events.
4.3.2 Experimental Results
Retraining Rank Correlation.
We form the set $U$ with 100 random subsets, each contains 50% of the training samples. We calculate the rank correlations on 178 generated music and report the average rank correlations for different data attribution methods in Table 1.
In comparison to the random baseline, both attribution methods have achieved significantly positive correlations with the ground-truth scores at the event level, and TRAK also works well at the segment level. This indicates that there exist computationally feasible solutions that can reasonably attribute the generated music to the copyrighted training content, thus solving the key technical bottleneck for establishing a royalty model. In addition, we observe that event-level attribution seems to be easier than segment-level attribution. This leads to an interesting question about the proper granularity of attributing generated music, which we leave for future exploration. For the rest of the paper, we conduct all the experiments with the TRAK-based attribution method.
Random TracIN TRAK
Segment-level 0.0091$\pm$0.007 -0.036$\pm$ 0.031 0.301$\pm$0.007
Event-level -0.0004$\pm$0.008 0.127$\pm$0.008 0.359$\pm$0.010
Table 1: The average retraining rank correlation among 178 generated music for different data attribution methods. The error bar represent the standard error of the mean. “Random” refers to a baseline that employs random attribution scores. “Segment-level” and “Event-level” refer to the two levels of attribution discussed in Section 4.2.1.
Musical Similarity.
For each generated music, we order and group the training music pieces by their attribution scores. Figure 1 shows the results of musical similarity correlation between the generated music and training music groups. We observe a clearly decreasing trend of musical similarity for training music group with lower attribution scores. This suggests that the data attribution methods also capture some influence in terms of musical styles (see Appendix E for detailed discussion).
Figure 1: Musical similarity in terms of loudness, key, and duration. The x-axis represents groups of training music pieces with decreasing TRAK attribution scores.
5 Robustness of Attribution Scores
In this section, we analyze the robustness of the TRAK attribution scores for music generative AI, which is crucial for establishing reliable royalty distribution. In Section 5.1, we examine the robustness against the randomness inherently existing in the data attribution process, which we term as stochastic robustness. In Section 5.2, we further investigate the adversarial robustness of the attribution scores against malicious actors that seek to adversarially manipulate the attribution scores.
5.1 Stochastic Robustness
Data attribution methods for deep neural networks naturally come with randomness due to, e.g., model initialization and training dynamics Søgaard et al. (2021); Nguyen et al. (2023).
In this subsection, we examine the stability of the attribution scores against such natural randomness. Specifically, we run the data attribution method multiple times with independent model initialization and training processes. Then we carry out Student’s t-test on the score of each training-generation pair (with the null hypothesis that the score equals to zero) as a way to quantify the stochastic robustness of the attribution scores.
(a) Histogram of the p-values for t-tests on attribution scores.
(b) P-value v.s. the ranking group of attribution scores.
Figure 2: P-values for t-tests on attribution scores.
Figure 2(a) shows the histogram of the p-values for t-tests on the training-generation pairs. We find that only a small portion of the pairs have p-values smaller than 0.05777Rigorously claiming statistical significance requires false-discovery-rate control for the multiple hypothesis testing. But we are only using the distribution of p-values to measure the stability of attribution scores, and do not intend to claim statistical significance for individual scores.. Furthermore, we group the training-generation pairs by the relative rankings of the average attribution scores for these pairs, and Figure 2(b) shows the boxplots of p-values for each group: the p-value is correlated with the rankings of the attribution scores. This result suggests that while the data attribution scores are not always stable, the ones with the top attribution scores tend to be reliable.
This result has an implication on the royalty mechanism design: the revenue of a generated music should be distributed to the training pieces with top attribution scores.
5.2 Adversarial Robustness
As with any system where financial interests are involved, there might be malicious actors who seek to increase their royalty shares by adversarially manipulating their contributed data. In this section, we evaluate the adversarial robustness of the TRAK attribution scores under two potential adversarial attack methods.
Duplication.
Duplicating a training sample multiple times is an intuitive way to increase its total attribution. In the experiment, we add multiple duplicate copies of a training sample into the dataset and then recalculate the attribution scores. Table 2 shows the total attribution score of the duplicated training samples over 178 generated music. We find that having duplications of a training sample in fact mostly decreases the total attribution score of this training sample and its duplications, indicating the attribution scores are robust against this type of attack (see Appendix F for explanations).
Copy-and-Replace.
Replacing part of one’s music with a segment from a highly influential training sample is another viable method to increase the attribution score of the altered music. In our experiment, we copy a segment from a training sample (source) that has top attribution scores and replace a segment of another training sample (target) with it. Table 3 presents how many times the modified target appears in the top-50 attribution scores among 178 generated music. It indicates that the attack can be effective even with a relatively small number of copied events.
The results of adversarial robustness demonstrate that the current best attribution method, TRAK, is robust to certain adversarial attacks. However, some attacks may still be successful if additional information, such as the attribution scores of other samples, is available. Enhancing the adversarial robustness of attribution scores represents a crucial direction for future research [^8] Please refer to Appendix F for detailed experimental setup..
Duplicate Copies 0 (original) 1 2 5 10 100
Sample 1 -4.187 0.410 -0.054 -0.007 0.115 0.014
Sample 2 0.188 -0.308 0.132 -0.004 -0.016 -0.009
Sample 3 5.128 -0.112 -0.058 -0.054 0.005 0.004
Table 2: The total attribution score of the duplicated sample over 178 generated music. Three training music sequences are chosen to be duplicated independently. Experiments with 1, 2, 5, 10, and 100 extra copies are conducted to compare against the baseline scenario of 0 copies, representing the original setting.
Copied Events 0 (original) 1 2 4 8 16 32
Source 0 0 17 25 24 41 23 33
Source 1 0 19 18 22 21 26 26
Table 3: The number of times that modified target appears in the top-50 attribution scores among 178 generated music (the higher the more influential). Two different music sequences with high attribution scores are selected independently as sources. Experiments on 1, 2, 4, 8, 16, and 32 copied events are conducted to compare against the baseline scenario of 0 events, representing the original setting.
6 Related Work
In this section, we discuss how the computational copyright solution and AI music royalty model in this paper are connected to prior works. These works can fall into three categories: (1) law, (2) economy, and (3) technology.
Law.
Legal issues related to generative AI copyright have attracted considerable interest. One of the primary discussions is the applicability of the fair use doctrine. Though the fair use doctrine was shown to be applicable in data mining through previous studies (Sag, 2018) and lawsuits [^9] See https://casetext.com/case/guild-v-google-inc-1, recent studies continue to raise questions about its suitability for AI-generated content (Henderson et al., 2023; Samuelson, 2023; Sag, 2023; Franceschelli and Musolesi, 2022; Peukert and Windisch, 2023). One of the main reasons for this transformation is that AI-generated content could be substantially similar to original works and impact the market (Henderson et al., 2023; Peukert and Windisch, 2023).
The applicability of fair use influences the determination of ownership and authorship regarding AI-generated content, an issue that remains actively debated and unresolved. The difficulty is to find a human authorship (Peukert and Windisch, 2023). Franceschelli and Musolesi (2022) argues that the author of AI generated works should be assigned to users. Lee et al. (2023) argues that the generation that does not meet fair use doctrine should be uncopyrightable unless further permission to derivative copyright.
Another challenge that is widely discussed is the test (Lee et al., 2023) to assert copyright infringement. The derivative works, like AI generation, may infringe on adaptation rights [^10] See Section 106 of the Copyright Act: https://www.copyright.gov/title17/92chap1.html#106., as long as proper tests are satisfied. It is suggested that the tests for copyright infringement can be divided into two main categories: access and similarity (Vyas et al., 2023; Lee et al., 2023). Access considers whether any part of an AI generation was directly influenced by a specific training example, and similarity considers if the generation is substantially similar to a training example. Correspondingly, we provide two perspectives, i.e., training data attribution and similarity, to define “influence” in Section 4.1 and show the potential synergy of these two perspectives. The legal community has recognized a technical gap in the process of determining which training samples influence the generation (Lim, 2023; Lee et al., 2023; Gans, 2024). This paper represents a pioneering effort aimed at bridging this gap.
Economy.
The effect of technical advancements, such as the digitalization of music, on the economic aspect of copyright has been discussed in numerous studies (Peukert and Windisch, 2023; Valdovinos Kaye et al., 2021; Gans, 2024). These discussions mainly focus on the social welfare of different copyright regimes. One of the most important regimes is Notice and Takedown (N&T) [^11] See https://www.copyright.gov/dmca/.. Systems following this regime are often employed by large platforms like YouTube (Van der Sar, 2021) and TikTok (Valdovinos Kaye et al., 2021). However, studies argue that these systems do not come without problems (Peukert and Windisch, 2023; Gans, 2015). Specifically, copyright owners, especially in the music industry (Gans, 2015), could benefit from derivative works like user-generated videos, but the forced taking down of derivative works undermines the welfare and promotion of copyright owners (Peukert and Windisch, 2023).
There is a trend that copyright owners are transitioning from using N&T systems to algorithmic licensing (Peukert and Windisch, 2023). Algorithmic licensing provides opportunities for creating a market where copyright owners could benefit from derivative works. For example, Content ID on YouTube provides monetizing option other than taking down for copyright owners. Empirical evidence indicates the monetizing option is widely chosen with a low dispute rate (cop, 2021). Remarkably, the royalty model and prototype proposed in Section 3 and Section 4 can be seen as an algorithmic licensing model for AI-generated music.
Technology.
There have been several studies of technical solutions in the copyright arena, while most of the existing literature focuses on preventing generative AI from generating infringing works (Vyas et al., 2023; Chu et al., 2023; Li et al., 2023; Somepalli et al., 2023b, a). Li et al. (2023); Somepalli et al. (2023b, a) investigate the memorization effect, which refers to the generation of something directly copied from training data, and the mitigations like deduplication in the diffusion model. Vyas et al. (2023); Chu et al. (2023) propose protection methods based on algorithm stability to prevent models from generating very similar works to copyrighted input. Though there have been promising results in this area, studies (Elkin-Koren et al., 2023) show that these methods could have some downsides that deviate from the purpose of copyright law.
Another closely relevant direction is data valuation, which usually assigns test sample agnostic value scores to training samples (Ghorbani and Zou, 2019). However, the model for music royalty necessitates a more fine-grained approach to attribution, extending to each test sample or, in more detailed instances, to each individual event.
7 Conclusion
Acknowledgement
The authors would like to thank Yi Liang for helpful discussions on evaluating musical styles and Hao Huang for some technical discussions.
Appendix A A Primer on the Concepts of Music Royalties